How many times more basic is sodium hydroxide (pH 14.0) than ammonia (pH 11.0)?
Answer #1

Solution :-

Using the given ph values lets calculate the concentration of the [OH-] for both

pOH= -log[OH-]

[OH-] = antilog [-pOH]

pOH for the NaOH = 0

therefore the concnetration of the NaOH that is OH^- = 1.0 M

and [OH-] from NH3 is calculated as

pOH = 14 - 11 = 3

[OH-] = antilog [-3]

[OH-] = 0.001 M

now lets calculate the ratio

NaOH / NH3 = 1.0 / 0.001 = 1000

Therefore NaOH is 1000 times greater base than NH3